# Adaptive Guardrails For Large Language Models via Trust Modeling and In-Context Learning

### v1

@@ Introduction
Additionally, Prompt Engineering is employed to measure area relevance ($AR$) by creating task-specific prompt templates that guide the inference process of LLMs ($LM(\cdot)\in[0,1]$). To further enhance the accuracy of the outcomes and reduce the impact of possible inference errors, we integrates normalized cosine similarity ($cosN(\cdot)\in[0,1]$) into our evaluation method. This quantitative assessment defined in Equation 12 allows the trust model to accurately evaluate the alignment between a trustee’s professional area demonstrated by TTPs and the sensitive prompts, ensuring that the system’s responses can meet diverse specialized needs:
$AR^{(\tau)}_{ki}=\max(cosN(Area,Input),LM(Area,Input))$ (12)
@@
Remark 1: We also incorporate periodic re-validation throughout the whole trust evaluation process, which is a common practice in web services (Ceccarelli et al. 2015), to ensure that all sensitive information requests meet specific security requirements within guardrail systems. This mandatory procedure applies universally, irrespective of trust scores, preventing outdated or overly trusted credentials from granting inappropriate access and thus enhancing the system’s robustness. For long-term access to sensitive or controversial information for legal or specialized purposes, users must periodically re-validate their identity with TTPs to maintain credential validity.
Contextual Adaptation by Guardrail Tiers
@@
Upon calculating user trust scores $T_{i}^{(\tau)}$, we propose classifying adaptive guardrails into different tiers that regulate the strictness of control mechanisms. This tiers enables precise management of users’ access level to sensitive information. Under this framework, key hyperparameters of LLMs such as temperature, maximum response length, and token limits can be dynamically adjusted based on the tier indicated by the user’s trust score, ensuring that the LLM operates within safe parameters while satisfying the personalization and demonstrated trustworthiness of different users. Additionally, ICL allows LLMs, governed by different guardrail tiers, to access hierarchical knowledge bases as each interaction context. This strategy ensures that the content depth and richness of the LLM’s responses are closely aligned with the constraints imposed by each guardrail tier.
The hierarchical structure of these knowledge bases facilitates a finely tuned response strategy. Users with higher trust score are granted access to more sensitive and confidential information. Given the trust score $T_{i}$ and LLM’s response $x$, The content richness function $R$ is defined below:
$R^{(\tau)}(x,T_{i})=LM\left(x,Context_{\min\left(\left\lfloor\frac{T_{i}^{(\tau)}}{\xi}\right\rfloor,L\right)}\right)$ (14)
where $LM(\cdot)$ generates content from the context tier depended on $\min\left(\left\lfloor\frac{T_{i}}{\xi}\right\rfloor,L\right)$, corresponding to the user’s calculated access level. Here, $\xi$ denote the access level threshold, and $L$ represents the total number of hierarchy of the knowledge base. Hence, users with lower trust levels engage with more generalized moderated content, restricted by an upper limit of accessible information layers.
By the adaptive trust modeling and contextual adaptation mechanism, our guardrail system significantly enhances the practicality of LLMs by ensuring their outputs align precisely with the ethical and safety standards required when sensitive information requested. Through dynamic tuning the model’s operational parameters and contextual knowledge base guided by overall trust score, the proposed guardrail mechanism establish a robust framework for safe and reliable LLM deployment across diverse user group, thus balancing the user expectations and content safety.
@@
Experiments
Data Preparation
For our experiments, we adapted AdvBench dataset (Zou et al. 2023), categorizing prompts by their relevance to computer science to evaluate our adaptive trust model for guardrails. This dataset comprises 520 harmful prompts reflecting potential sensitive interactions, supplemented by 3,000 safe prompts for interaction diversity. We specifically selected 162 harmful prompts highly relevant to CS and 358 from other domains to test the guardrail’s efficacy in managing access to sensitive information. Due to privacy concerns and restricted access to actual user data from TTPs, we created synthetic profiles for virtual users, incorporating simulated ratings, professional areas, social information, and more, to emulate real-world trust attributes, allowing us to verify if the trust model effectively aligns access to sensitive information with validated user identities. All necessary materials including code and datasets, are hosted on GitHub [^8] The anonymized GitHub will be made public upon paper acceptance to comply with double-blind review requirements.
Implementation Assumptions
@@
Case 1: Adaptability and Security Analysis
@@
In comparison with black-box techniques such as the Llama Guard series and Nvidia NeMo, our system exhibited enhanced control and safety. Llama Guard allowed 96.30% access to CS-related content, slightly lower than our method. This discrepancy may stem from pre-trained biases, such as strongly associating words like “virus” with medical contexts, reflecting the necessity for fine-tuning in dynamic scenarios of existing guardrails. Llama Guard also allowed 89.94% access to unrelated areas, highlighting a lack of precision that could introduce security risks. On the other hand, Llama Guard 2 and 3 significantly restricted access, potentially to the point of impeding user experience due to over-restrictiveness. Nvidia NeMo, with more moderate access rates of 48.15% and 40.94% to CS and other areas respectively, suggested a balanced approach but does not match the nuanced control our model provides. Additionally, the results revealed that black-box methods predominantly employed rigid, predefined rules. Despite our efforts to slightly relaxed these to accommodate the specialized needs of trustworthy users, they still had similar probabilities, indicating the difficulty in differentiating sensitive queries. While Llama Guard and Nvidia NeMo enable customization of security settings for specific queries, they demonstrate vulnerabilities in preventing unauthorized access to sensitive content. Our system addresses these shortcomings by implementing access controls based on verifiable user credentials and query relevance. Our approach secures sensitive information and adheres to ethical standards, establishing our adaptive guardrail system as a pioneering solution for the secure and flexible deployment of future LLMs.
@@
Case 2: Contextual Implications Analysis
Case 3: Ablation Test Analysis
@@
Conclusion
This paper presents a groundbreaking approach to safeguarding the security and utility of large language models through an adaptive guardrail mechanism based on trust modeling and In-context Learning. Our model dynamically integrates real-time, user-specific trust assessments, offering personalized content moderation previously unachievable with static guardrail technologies. This trust-oriented method effectively secures sensitive information without compromising user engagement, thereby meeting diverse user needs. Future work could be extended to include more comprehensive contextual hierarchy and deeper integration with emerging AI technologies, refining the dimension of trust-oriented assessments in complex interaction scenarios. This foundational work paves the way for responsible AI advancements, promoting more dependable and user-centric deployment of AI models.

### v2

@@ 3.0.2 Authority-Verified Trust
To enhance the precision of area relevance ($AR$) assessments within our trust model, we employ Prompt Engineering to direct the inference process of multiple LLMs debating $\mathcal{D}$. The robustness of our final decision is strengthened by integrating semantic similarity measures, denoted as $ST$. Specifically, we employ the all-mpnet-base-v2 sentence transformer in our cases, which yields semantically rich embeddings that effectively capture nuanced linguistic relations Reimers and Gurevych (2019). This method refines the evaluation of area relevance between a trustee’s professional area, as verified by TTPs, and the sensitivity required by the prompts. The formula for computing $AR$ in $[0,1]$ is presented as follows:
$AR^{(\tau)}_{ki}=\frac{ST^{(\tau)}(Area,Input)+\mathcal{D}^{(\tau)}(Area,Input)}{2}$ (12)
@@
In this paper, we consider a set of TTPs $p_{k}\in P$, where $k$ is the TTP index and these TTPs are divided by three ranking levels $\mathcal{A}.ranking\in\{top,medium,low\}$. Each level is assigned a weight indicator $\mathcal{A}\in[0,1]$, which quantifies their relative authority. Furthermore, TTPs are independent entities that provide verification services for systems, ensuring the authenticity and accuracy of data without direct involvement in the transactions or interactions between primary parties. For instance, validating a university email to confirm educational qualifications illustrates the necessity of TTPs in accessing specialized information for research purposes.
@@
Remark 1: We also incorporate periodic re-validation throughout the whole trust evaluation process, which is a common practice in web services Ceccarelli et al. (2015), to ensure that all sensitive information requests meet specific security requirements within guardrail systems. This mandatory procedure applies universally, irrespective of trust scores, preventing outdated or overly trusted credentials from granting inappropriate access and thus enhancing the system’s robustness. For long-term access to sensitive or controversial information for legal or specialized purposes, users must periodically re-validate their identity with TTPs to maintain credential validity.
3.1 Contextual Adaptation Mechanism
@@
Upon calculating user trust scores, we classify adaptive guardrails into tiers that regulate control mechanisms. These tiers manage users’ access to sensitive information. Key hyperparameters of LLMs, such as temperature, response length, and token limits, are dynamically adjusted based on the user’s trust score, ensuring safe operation while accommodating personalization and trust. For example, users with higher trust scores may have higher token limits for generating responses when accessing sensitive information.
Figure 3: Contextual Adaptation via RAG and ICL
Furthermore, integrating RAG and ICL significantly enhances the relevance and specificity of responses in language models. In our system, RAG leverages an expertly constructed, hierarchically structured confidential knowledge base to retrieve trust-oriented contexts. This integration ensures that the response generation process is dynamically tailored according to both the contexts and the user’s trust score. The confidentiality of the access level is governed by $T_{i}$ and is defined as follows:
$AL_{i}=\sum_{l=1}^{L}\mathbf{1}(T_{i}\geq\xi_{l})$ (14)
where $\xi_{l}$ represents the threshold for the $l$-th access level, and $L$ denotes the total number of hierarchical layers within the knowledge base. This formula employs an indicator function $\mathbf{1}(\cdot)$, which assigns a value of 1 for each threshold that the trust score $T_{i}$ meets or exceeds, thereby cumulating the number of accessible layers. This structure allows for fine-tuned access control, where higher trust scores permit access to more sensitive information, and lower scores restrict users to more generalized content. The process is depicted in Figure 3, illustrating the integration of RAG and ICL to manage information flow based on the user’s trust score. Initially, RAG retrieves trust-oriented contexts, which are then seamlessly integrated by ICL along with a guidance prompt and the trust score, to guide the response generated by the language model. This tailored approach ensures a balance between content safety and personalization, effectively addressing specific user needs. The content richness function is therefore formulated as below:
$R^{(\tau)}_{i}(x,AL_{i})=LM\left(x,Context_{AL_{i}}\right)$ (15)
In this formula, $LM(\cdot)$ tailors response $x$ according to the user’s authorized access level $AL_{i}$, specifically to the $Context$ that pertains to the query. Here, $Context$ refers to the enriched retrieval information relevant to the inquiry. This adaptive guardrail mechanism ensures each user interaction is secure and personalized, balancing usability with security protocols.
By employing adaptive trust modeling along with content adaptation mechanism, our guardrail not only enhances the practicality of LLMs but also ensures outputs adhere to ethical and safety standards. This dynamic adjustment of model parameters and sensitive knowledge access level based on trust scores establishes a framework for the safe and reliable deployment of language models, effectively balancing user expectations with content integrity.
@@
4 Experiments
4.1 Data Preparation
For our experiments, we adapted AdvBench dataset Zou et al. (2023), categorizing prompts by their relevance to computer science (CS) to evaluate our adaptive trust model for guardrails. This dataset comprises 520 harmful prompts reflecting potential sensitive interactions, supplemented by 3,000 safe prompts for interaction diversity. We specifically selected 162 harmful prompts highly relevant to CS and 358 from other domains to test the guardrail’s efficacy in managing access to sensitive information. Due to privacy concerns and restricted access to actual user data from TTPs, we created synthetic profiles for virtual users, incorporating simulated ratings, professional areas, social information, and more, to emulate real-world trust attributes, allowing us to verify if the trust model effectively aligns access to sensitive information with validated user identities. All necessary materials including code and datasets, are hosted on GitHub [^4] The anonymized GitHub will be made public upon paper acceptance to comply with double-blind review requirements.
4.2 Implementation Assumptions
@@
4.3 Case 1: Adaptability and Security Analysis
@@
In comparison with black-box techniques such as the Llama Guard series and Nvidia NeMo, our system exhibited enhanced control and safety. Llama Guard allowed 96.30% access to CS-related content, slightly lower than our method. This discrepancy may stem from pre-trained biases, such as strongly associating words like “virus” with medical contexts, reflecting the necessity for fine-tuning in dynamic scenarios of existing guardrails. Llama Guard also allowed 89.94% access to unrelated areas, highlighting a lack of precision that could introduce security risks. On the other hand, Llama Guard 2 and 3 significantly restricted access, potentially to the point of impeding user experience due to over-restrictiveness. Nvidia NeMo, with more moderate access rates of 48.15% and 40.94% to CS and other areas respectively, suggested a balanced approach but does not match the nuanced control our model provides. Additionally, the results revealed that black-box methods predominantly employed rigid, predefined rules. Despite our efforts to slightly relaxed these to accommodate the specialized needs of trustworthy users, they still had similar probabilities, indicating the difficulty in differentiating sensitive queries. While Llama Guard and Nvidia NeMo enable customization of security settings for specific queries, they demonstrate vulnerabilities in preventing unauthorized access to sensitive content. Our system overcomes these shortcomings by implementing access controls through verifiable user credentials and query relevance. This approach safeguards sensitive information while upholding ethical standards, positioning our adaptive guardrail system as a groundbreaking solution for the secure and versatile deployment of future LLMs. Details of the ablation study for each variables are provided in Appendix B.
5 Case 2: Contextual Implications Analysis
Figure 5: Trust Dimension Contributions to Dynamic Trust Modeling Across Authority Levels
5.1 Case 3: Trust Dimension Contribution Across Varying Authority Levels
To further illustrate the adaptability and rationality of our trust computation framework across users authorized by TTPs with varying authority levels, we conducted a trust contribution demonstration of how trust dynamics facilitate or restrict access to sensitive information, as visualized in Figure 5. This demonstration assessed users with similar basic trust attributes, while their authority levels determined the respective contributions of $DT$ and $AT$ to the final $T_{i}$.
For users with Top Authority credentials, $T_{i}$ was predominantly influenced by $AT$, given that their access rights were validated by highly reliable TTPs. The dominance of $AT$ ensured that such users could securely and stably access sensitive content within their authorized domains. Moreover, this design maintained robust safeguards to prevent misuse by aligning access privileges with the verifiable reliability of their top authority credentials. In contrast, users without Top Authority credentials were subject to a more dynamic combination of $DT$ and $AT$. For these users, $DT$ derived from their historical interactions played a pivotal role, especially in moderating ongoing access. Even when their $AT$ was moderately high, reflecting some degree of validation by less authoritative TTPs, their $DT$ scores provided the necessary flexibility to adapt access levels in real time. This combination ensured that users demonstrating consistent, trustworthy behavior were allowed controlled access to relevant sensitive information, while prolonged or excessive engagement with such content was restricted to mitigate potential risks. The system dynamically adjusted the weight of $DT$ to compensate for the lack of strong $AT$ validation, enhancing its capacity to adapt to diverse trust scenarios.
For users with low $DT$ and low authority levels, the framework enforced strict restrictions, effectively blocking access to sensitive information. This was true even when less credible TTPs assigned these users seemingly high trust scores, as the system integrated multi-dimensional trust attributes to prevent exploitation of falsified credentials. By combining both $DT$ and $AT$, the model demonstrated robustness against potential vulnerabilities, ensuring that unauthorized users could not bypass the guardrails. This design significantly enhances the system’s reliability, particularly in scenarios involving complex or adversarial interactions.
The experimental results highlight the distinct contributions of $DT$ and $AT$ to trust computation, with their influence varying by authority level. For Top Authority users, $AT$ dominated due to highly credible credentials, while $DT$ played a pivotal role for lower authority users by leveraging historical interactions to refine trust scores dynamically. This interplay underscores the framework’s adaptability in balancing security, usability, and flexibility. By integrating direct and authority-verified trust, the model enforces robust safeguards while enabling context-aware adjustments for secure, personalized access. Appendix C further demonstrates its enhanced resilience against jailbreaking attacks compared to static LLM-based guardrails.
6 Conclusion
This paper presents a groundbreaking approach to safeguarding the security and utility of large language models through an adaptive guardrail mechanism based on trust modeling and online ICL via RAG. Our model dynamically integrates real-time, user-specific trust assessments, offering personalized content moderation previously unachievable with static guardrail technologies. This trust-oriented method effectively secures sensitive information without compromising user engagement, thereby meeting diverse user needs. Future work could be extended to include more comprehensive contextual hierarchy and deeper integration with emerging AI technologies, refining the dimension of trust-oriented assessments in complex interaction scenarios. This foundational work paves the way for responsible AI advancements, promoting more dependable and user-centric deployment of AI models.
Appendix
@@ 6 Conclusion
Appendix A Related Work
A.1 Existing Guardrails
Guardrails for LLMs are programmable, rule-based systems acting as intermediaries to ensure that interactions between users and models adhere to established ethical and operational standards Dong et al. [2024b]. Reflecting the latest advancement, several existing techniques represent the state of the art in this field. Llama Guard Inan et al. [2023] operates as a binary classifier that effectively differentiates between safe and unsafe content, ensuring moderation aligns with predefined safety guidelines. Concurrently, Nvidia NeMo Rebedea et al. [2023] enhances interaction accuracy by utilizing the K-nearest neighbors (KNN) method to align user intents with the most appropriate vector-based canonical forms. After this matching process, LLMs generate safe and pertinent responses, guided by Colang scripts. Moreover, Guardrails AI Rajpal [2023] bolsters system reliability and trust by implementing structured safety protocols, involving detailed XML specifications that rigorously check that outputs adhere to strict content integrity and compliance standards. Supporting these technologies, Python packages such as LangChain [^5] https://www.langchain.com/, AI Fairness 360666https://github.com/Trusted-AI/AIF360, Adversarial Robustness Toolbox [^7] https://github.com/Trusted-AI/adversarial-robustness-toolbox, Fairlearn [^8] https://github.com/fairlearn/fairlearn, and Detoxify [^9] https://github.com/unitaryai/detoxify address biases, boost robustness, and ensure content safety and fairness, collectively fostering responsible and reliable LLM applications.
A.2 Trust Models
Trust is defined by the Merriam-Webster Dictionary Dictionary [2024] as “assured reliance on the character, ability, strength, or truth of someone or something,” which succinctly encapsulates trust as a fundamental aspect of relationships, where a trustor places confidence in a trustee based on established criteria. Recognized as a pivotal element in decision-making, trust spans various disciplines including management, psychology, economics, and engineering, underscoring its multidisciplinary importance Lahijanian and Kwiatkowska [2016].
Recently, trust in human-technology relationships has gained widespread attention and is typically classified into three distinct types: credentials-based, cognitive trust, and experience-based Huang et al. [2019]. Credentials-based trust is commonly employed in security systems, where users are required to provide valid credentials to verify their identity and gain access to services, satisfying established security policies Marsh and Dibben [2003]. Cognitive trust involves in the human aspects of trust, focusing on subjective judgments and trust-based decisions in interactions, especially between humans and robots, highlighting the psychological dimensions of trust Falcone and Castelfranchi [2001]. Experience-based trust is frequently used in multiagent systems and e-service platforms, where a trustee’s reliability is assessed based on past interactions and reputation-based trust, helping to predict the trustees’ future behaviors Griffiths [2005]; Şensoy et al. [2007]. This approach helps predict future behaviors by employing statistical methods to calculate trust scores. In this paper, we introduce a composite trust model that identifies relevant dimensions from various trust paradigms to meet the dual needs of security and practicality in LLM-user interactions, comprehensively assessing trustee reliability in LLMs equipped with adaptive guardrails.
A.3 In-Context Learning
In-context learning utilizes the generalization capabilities of LLMs to efficiently perform tasks with just a few contextual examples, requiring minimal data and no parameter updates, setting it apart from standard in-weights learning like gradient-based fine-tuning Kossen et al. [2024]. This method has become a focal point due to its ability to streamline the learning process and reduce the computational overhead typically associated with training large models. Research has focused on the selection of demonstration examples and the better ways of problem formulations Zhao et al. [2021]; Liu et al. [2021]. A common enhancement to ICL is RAG, which retrieves external knowledge to refine LLM outputs with contextually rich information, improving accuracy and adaptability in real-world scenarios Singal et al. [2024]; Fan et al. [2024]. In our paper, we extend this approach by leveraging RAG to implement online in-context learning, dynamically retrieving trust-oriented information from hierarchical knowledge bases. By incorporating trust-based criteria, our method ensures that the retrieved knowledge is not only contextually rich but also aligned with user-specific constraints, enabling LLMs to interact effectively with external agents or access expert-curated knowledge sources.
Appendix B Ablation Test Analysis
@@
Appendix C Trust-Driven Safeguards in Mitigating Jailbreaking Threats
The trust-oriented adaptive guardrail system establishes a more robust safeguards against jailbreaking attacks by leveraging a dual-layered trust computation framework comprising $DT$ and $AT$. Unlike static guardrails, which are prone to being bypassed through carefully crafted prompts, this adaptive system preemptively assesses the user’s trust level before applying normal guardrail logic to determine whether a request involving sensitive information should be accepted or denied. This design significantly enhances the robustness of the system, particularly against adversarial scenarios involving jailbreaking attempts.
Attackers often exploit static guardrails by submitting a mix of benign and sensitive queries to mask their malicious intent. However, the adaptive guardrail counters such strategies by dynamically calibrating $DT$ and $AT$ scores based on historical interaction data and authority validation. For instance, attackers with limited interaction histories exhibit low $DT$, as their lack of prior trustworthy engagement raises suspicion. Similarly, users failing to demonstrate validation from reliable TTPs are assigned low $AT$, further restricting their access. This trust-oriented approach ensures that users with insufficient trust scores are denied access outright, regardless of the crafted nature of their prompts.
To illustrate this capability, Figure 6 presents a simulation involving a low-authority user attempting to bypass the guardrail system using mixed benign and sensitive jailbreaking queries. Unlike high-authority users, whose trust scores predominantly rely on TTP-based validation and are robust against malicious attack strategies, low-authority users’ trust scores are dynamically adjusted by $DT$. Sensitive queries cause a cumulative decline in $DT$, progressively eroding the user’s trust score. This mechanism effectively denies access, even when intermittent safe prompts are submitted to obscure malicious intent. Additionally, periodic $DT$ resets and stricter penalty constraints prevent attackers from exploiting the system by repeatedly submitting benign queries to regain access.
Figure 6: Mitigating Jailbreaking Attacks Through Direct Trust Decline from Sensitive Queries
The adaptive guardrail’s robustness is rooted in its trust modeling foundation, which integrates historical user data and authority validation to preemptively assess trust levels before applying content moderation logic. By ensuring only users meeting trust thresholds can proceed, the system mitigates unauthorized access attempts early in the process. Unlike static guardrails, which lack adaptability against dynamic attack patterns, this layered approach dynamically identifies and blocks adversarial users through continuous trust monitoring and recalibration. Users attempting to exploit the system with sophisticated prompts or mixed benign and sensitive queries are quickly flagged due to their low trust profiles, effectively preventing prolonged exploitation. This trust-oriented mechanism also ensures that sensitive content is accessible only to high-trust users while maintaining robust protections against unauthorized access and misuse. By balancing personalization, usability and security, the system demonstrates resilience against jailbreaking attacks and establishes a scalable framework for trust-aware LLM applications in security-critical environments.